a+b=6
b+c=5
c+d=4
d+e=2
e+a=2
Solve for all.

Answer by boilpoil(127)

if it is like this then we should 'combine' the equations all together:

a+b+b+c+c+d+d+e+e+a=6+5+4+2+2 ---Adding all questions together is the same as adding all the answers together
2a+2b+2c+2d+2e=19
2(a+b+c+d+e)=19
a+b+c+d+e=9.5
a+6-a+c+4-c+2-a=9.5 ---By carefully choosing which variable to substitute in, we are able to make this a simple equation
6+4+2-a=9.5
12-a=9.5
-a=-2.5
a=2.5

b=6-2.5
=3.5

c=5-3.5
=1.5

d=4-1.5
=2.5

e=2-2.5
=-0.5

Doing these types of questions requires a flexible thinking, you might need to train yourself on these types of questions more.